# Automatic Input Rewriting Improves Translation with Large Language Models

### v1

@@ 2 Input Rewriting Methods
2.1 MT-Agnostic Rewriting
@@
2.2 Task-Aware Rewriting
@@
2.3 Translatability-Aware Rewriting
@@ 3 Experimental Setup
MT-Agnostic Simplification (Tower) 0.915 0.907 1.504* 1.519
@@
Task-Aware Easy translate (Tower) 0.901 0.903 1.759 2.427
@@
Translatability-Aware Selection 0.921* 0.922* 1.734 1.461*
@@
MT-Agnostic Simplification (Tower) 0.921* 0.891 1.135 1.921
@@
Task-Aware Easy translate (LLaMA-3) 0.917 0.881 0.801* 10.401
@@
Translatability-Aware Selection 0.914 0.899* 2.096 1.830*
@@
MT-Agnostic Simplification (Tower) 0.821 0.802 1.521* 2.227
@@
Task-Aware Easy translate (LLaMA-3) 0.793 0.791 1.618 7.650
@@
Translatability-Aware Selection 0.823* 0.819* 3.149 2.206*
@@ Appendix F Time & Computational Efficiency
MT-Agnostic Simplification (LLaMA-2) 0.931 0.846 0.900 1.185 1.727
@@
Task-Aware Easy Translation (LLaMA-2) 0.916 0.857 0.893 1.654 2.482
@@
Translatability-Aware Selection 0.921 0.907 0.915* 1.734 1.461*
@@
MT-Agnostic Simplification (LLaMA-2) 0.916 0.839 0.882 0.951 2.160
@@
Task-Aware Easy Translation (LLaMA-2) 0.914 0.839 0.884 1.037 10.849
@@
Translatability-Aware Selection 0.914 0.891* 0.899* 2.096 1.830*
@@
MT-Agnostic Simplification (LLaMA-2) 0.828 0.728 0.796 1.321 2.537
@@
Task-Aware Easy Translation (LLaMA-2) 0.821 0.721 0.784 1.900 7.732
@@
Translatability-Aware Selection 0.823 0.783* 0.819* 3.149 2.206*

### v2

@@ 2 Input Rewriting Methods
2.1 [Uncaptioned image] MT-Agnostic Rewriting
@@
2.2 [Uncaptioned image] Task-Aware Rewriting
@@
2.3 [Uncaptioned image] Translatability-Aware Rewriting
@@ 3 Experimental Setup
[Uncaptioned image] MT-Agnostic Simplification (Tower) 0.915 0.907 1.504* 1.519
@@
[Uncaptioned image] Task-Aware Easy translate (Tower) 0.901 0.903 1.759 2.427
@@
[Uncaptioned image] Translatability-Aware Selection 0.921* 0.922* 1.734 1.461*
@@
[Uncaptioned image] MT-Agnostic Simplification (Tower) 0.921* 0.891 1.135 1.921
@@
[Uncaptioned image] Task-Aware Easy translate (LLaMA-3) 0.917 0.881 0.801* 10.401
@@
[Uncaptioned image] Translatability-Aware Selection 0.914 0.899* 2.096 1.830*
@@
[Uncaptioned image] MT-Agnostic Simplification (Tower) 0.821 0.802 1.521* 2.227
@@
[Uncaptioned image] Task-Aware Easy translate (LLaMA-3) 0.793 0.791 1.618 7.650
@@
[Uncaptioned image] Translatability-Aware Selection 0.823* 0.819* 3.149 2.206*
@@ Appendix F Time & Computational Efficiency
[Uncaptioned image] MT-Agnostic Simplification (LLaMA-2) 0.931 0.846 0.900 1.185 1.727
@@
[Uncaptioned image] Task-Aware Easy Translation (LLaMA-2) 0.916 0.857 0.893 1.654 2.482
@@
[Uncaptioned image] Translatability-Aware Selection 0.921 0.907 0.915* 1.734 1.461*
@@
[Uncaptioned image] MT-Agnostic Simplification (LLaMA-2) 0.916 0.839 0.882 0.951 2.160
@@
[Uncaptioned image] Task-Aware Easy Translation (LLaMA-2) 0.914 0.839 0.884 1.037 10.849
@@
[Uncaptioned image] Translatability-Aware Selection 0.914 0.891* 0.899* 2.096 1.830*
@@
[Uncaptioned image] MT-Agnostic Simplification (LLaMA-2) 0.828 0.728 0.796 1.321 2.537
@@
[Uncaptioned image] Task-Aware Easy Translation (LLaMA-2) 0.821 0.721 0.784 1.900 7.732
@@
[Uncaptioned image] Translatability-Aware Selection 0.823 0.783* 0.819* 3.149 2.206*